OpenStudy (anonymous):

Find the inverse function for f(x)= (ln√x)/5

terenzreignz (terenzreignz):

Okay, so you have \[\large y = \frac{\ln\sqrt x}{5}\] right? What you wanna do is switch y and x first, like this... \[\large x = \frac{\ln\sqrt y}{5}\] and see if you can get y to stand alone; all by itself.